A health shop ownder made trail mix containing dried fruit, nuts, and carob chips. The dried fruit sells for $5.50 per pound, the nuts sell for $7.50 per pound, and the carob chips sell for $8.50 per pound. The shop owner mixed 50 pounds of the trail mix and sells it for $6.70 per pound. If the amount of nuts is five pounds more than the amount of carob chips, how much of each item was used for the trail mix?
Answer by jorel1380(3719)

f+n+c=50
c+5=n
50x6.70=335
f+c+5+c=50
f+2c=45
550f+750(c+5)+850c=33500
550f+750c+3750+850c=33500
550f+1600c=29750
550f+1100c=24750
500c=5000
c=10
n=15
f=25
The health shop owner needs 25 lbs. of fruit, 15 lbs. of nuts, and 10 lbs. of carobs for his mix..